# Understanding the Uncertainty of LLM Explanations: A Perspective Based on Reasoning Topology

### v1

@@
Abstract.
Understanding the uncertainty in large language model (LLM) explanations is important for evaluating their faithfulness and reasoning consistency, and thus provides insights into the reliability of LLM’s output regarding a question. In this work, we propose a novel framework that quantifies uncertainty in LLM explanations through a reasoning topology perspective. By designing a structural elicitation strategy, we guide the LLMs to frame the explanations of an answer into a graph topology. This process decomposes the explanations into the knowledge related sub-questions and topology-based reasoning structures, which allows us to quantify uncertainty not only at the semantic level but also from the reasoning path. It further brings convenience to assess knowledge redundancy and provide interpretable insights into the reasoning process. Our method offers a systematic way to interpret the LLM reasoning, analyze limitations, and provide guidance for enhancing robustness and faithfulness. This work pioneers the use of graph-structured uncertainty measurement in LLM explanations and demonstrates the potential of topology-based quantification. The response data and code will be released upon publication.
Uncertainty Quantification, Natural Language Explanation, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Uncertainty quantificationccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Explainable AI
1. Introduction
@@
Deep learning models have long been criticized for their lack of trustworthiness due to their complex network structures and opaque decision-making processes (Li et al., 2022; Doshi-Velez and Kim, 2017; Samek et al., 2021). This has motivated researchers to investigate methods for understanding and quantifying the uncertainty associated with these models (Abdar et al., 2021; Loquercio et al., 2020; Maddox et al., 2019). Recently, Large Language Models (LLMs) have demonstrated significant advancements over traditional deep learning approaches across a variety of tasks (Zhao et al., 2023; Naveed et al., 2023). However, concerns about their reliability persist. LLMs often produce outputs that are difficult to verify, particularly in scenarios requiring complex reasoning (Shi et al., 2024). This introduces risks in critical applications, such as healthcare or legal domains (Cascella et al., 2023; Jayakumar et al., 2023), where incorrect or unreliable reasoning can have severe consequences. Properly quantifying uncertainty in the reasoning processes of LLMs is therefore crucial for ensuring their safe and effective deployment.
Existing research on Uncertainty Quantification (UQ) for LLMs primarily focuses on analyzing semantic uncertainty (Kuhn et al., 2023; Lin et al., 2023; Qiu and Miikkulainen, 2024; Da et al., 2024), which involves examining patterns in the meaning and phrasing of multiple responses generated for a given question. Although this approach provides insights into output-level variability, it neglects the logical reasoning steps that lead to these answers. As a result, it fails to address foundational issues in the reasoning process that could be used to debug or optimize model outputs. For instance, when asked by the the same question, different reasoning paths may converge on the same final answer (an example shown in Appendix Figure 6), yet some paths may involve inconsistencies or logical leaps. By quantifying uncertainty at the level of reasoning steps, we can better identify such inconsistencies, support human-in-the-loop systems for validating outputs in sensitive applications, and uncover weaknesses in a model’s reasoning process. This highlights the importance of incorporating reasoning-based uncertainty quantification.
In this paper, we address the problem of uncertainty quantification for logical reasoning steps by explicitly modeling reasoning processes as logical topologies. Existing work often treats the reasoning process generated by single Chain-of-Thought (CoT) sequences as a single ”long answer” (Wei et al., 2022; Wang et al., 2023), calculating semantic consistencies directly. While this approach captures some aspects of reasoning paths, it oversimplifies real-world reasoning processes, which often involve hierarchical dependencies and parallel sub-tasks. To overcome these limitations, we propose a novel formalism that explicitly models reasoning as a logical graph. In this representation, nodes correspond to individual reasoning steps, while edges capture logical dependencies between them. This structure enables more granular and interpretable analyses of uncertainty.
@@
  • We identify limitations in existing UQ approaches for LLMs and propose a novel framework that explicitly models the reasoning process from explanations as a logical topology.
  • We introduce multiple measures, including graph-based uncertainty measures and redundancy metrics, to provide granular reasoning variance and interpretable assessments of redundancy.
  • We demonstrate the effectiveness of our framework through extensive experiments across multiple datasets and LLMs. Our results highlight its ability to identify inconsistencies in reasoning paths, improving trustworthiness, and interpretability. Additionally, we discover three widely adopted reasoning patterns in LLMs and show the chance of improvement under the redundancy measure.
2. Related Work
@@ 2.1. UQ for LLM
A significant body of research has focused on performing UQ for LLMs by inducing the models to output their uncertainty along with their responses (Kadavath et al., 2022; Lin et al., 2022; Mielke et al., 2020; Tian et al., 2023). These methods often rely on token-level probabilities to train or fine-tune models for predicting uncertainty. While effective, these approaches require full access to the model’s structure and weights, which is impractical for black-box or commercial LLMs. For example, supervised methods such as those in (Kadavath et al., 2022) estimate uncertainty using logits and ground truth labels but are computationally expensive and resource-intensive.
@@
Another line of work estimates uncertainty directly at the response level using semantic entropy (Kuhn et al., 2023). While this method avoids token-level dependencies, it still relies on access to token probabilities, limiting its applicability in black-box settings. To address these limitations, researchers have proposed lightweight black-box methods that analyze response inconsistencies. For instance,  Lin et al. (2023) use graph Laplacian eigenvalues as an uncertainty indicator, while  Chen and Mueller (2023) computes confidence scores from generated outputs to identify speculative or unreliable answers. However, these approaches primarily focus on semantic-level analysis and neglect the logical structure underlying reasoning processes. Moreover, methods like (Lin et al., 2023) average entailment probabilities without considering directional information in reasoning paths.
@@
Our work is agnostic to the white box or black box since it leverages the generated explanations as a proxy to measure the reasoning uncertainty as in Figure 1. This enables a more nuanced and interpretable assessment of uncertainty in reasoning processes.
2.2. UQ for LLM Explanation
@@
Explanation-based UQ focuses on assessing the reliability of natural language explanations (NLEs) generated by LLMs by either prompting models to express confidence in their explanations or analyzing consistency across multiple outputs under varying conditions (Tanneru et al., 2024; Yadkori et al., 2024). While these methods provide insights into explanation robustness, they treat explanations to a question as a unstructured text representation, which lacks structural information and fails to capture inconsistencies or leaps in logic. In contrast, our work explicitly leverages well-structured reasoning topologies to enhance the UQ process for explanations. This structured representation enables us to assess explanation uncertainties at a finer granularity within complex reasoning paths.
3. Preliminaries
@@
In this section, we provide the foundational concepts for our study, including the definition of uncertainty in LLMs and the quantification of uncertainty in natural language explanations (NLE).
3.1. Uncertainty of LLMs
Uncertainty quantification (UQ) has been a critical topic in classical machine learning tasks such as classification and regression (Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Hernández-Lobato and Adams, 2015; Abdar et al., 2021). However, natural language generation (NLG) poses unique challenges for uncertainty quantification due to its high-dimensional output space, the need to account for semantic equivalence across distinct token sequences, and the limited accessibility of internal model outputs in black-box LLMs (Lin et al., 2023).
Figure 1. The category of the LLM uncertainty measure. It can be divided into two types, direct measure and proxy-based measure (use the explanations of each answer).
Given these challenges, most UQ methods for LLMs focus on analyzing uncertainty directly from model-generated responses as shown in the left part of Figure. 1. The problem could be defined as follows:
@@ 3.1. Uncertainty of LLMs
When an LLM $M$ is provided with an input $x$, either a query or prompt, the goal of an uncertainty function $U_{x}$ is to map the generated outputs to a scalar score that determines the uncertainty of the LLM $M$, i.e.,
(1) $U_{x}=\mathcal{U}\left(\{M(x_{i})\}_{i=1}^{n}\right)$
Here, $\{M(x_{i})\}_{i=1}^{n}$ denotes a set of $n$ responses generated by the model $M$, and $\mathcal{U}$ aggregates uncertainty across multiple responses. Note that while noticed that some work models have uncertainty describing the confidence of a specific output given the input (Tanneru et al., 2024), in this paper $U_{x}$ only depends on $x$ and is a property of the predicted distribution, which is estimated by $\mathcal{U}$ that aggregates uncertainty across multiple responses.
Depending on the specific UQ method used, each $x_{i}$ may correspond to either repeated inputs or rephrased variations of the input prompt $x$. For black-box methods that analyze response variability or semantic consistency (Lin et al., 2023), multiple outputs ($n>1$) are typically required. In contrast, white-box methods that rely on internal model information such as logits (Kadavath et al., 2022) may only require a single output ($n=1$). However, both black-box and white-box methods have limitations when applied to reasoning tasks. Black-box methods often focus on semantic output variability without capturing deeper uncertainties in reasoning steps. White-box methods are frequently inaccessible due to API restrictions or computational constraints.
3.2. Uncertainty of LLM Explanations
To address these issues, researchers try to understand uncertainties in reasoning processes through the natural language explanations (NLEs) (Camburu et al., 2018; Tanneru et al., 2024) as a proxy. As shown in the right part of Figure. 1, An NLE is a textual reasoning sequence generated by a language model $M$, typically derived to justify or explain the answer $a$ for a given input question $x^{q}$. We formally define it as follows:
@@ 3.2. Uncertainty of LLM Explanations
Given a model $M$ and an input prompt $x^{q}$, an NLE can be represented as:
(2) $M(x^{q}+x^{e})=a+a^{e}$
where $x^{e}$ is an explanation-specific prompt, $a$ is the model’s answer to the query $x^{q}$, and $a^{e}$ is the generated explanation accompanying the answer.
The explanation $a^{e}$ contains a sequence of reasoning steps, represented as $a^{e}=\{s_{1},s_{2},...,s_{m}\}$, which capture the reasoning process or justification for $a$.
To quantify uncertainty in explanations, we extend Problem 1 to consider $n$ explanations generated for the same query $x^{q}$. Each explanation $a^{e}_{i}$ ($i\in\{1,2,\dots,n\}$) corresponds to a set of reasoning steps derived from the same query. Each explanation consists of $m$ reasoning steps, represented as $a^{e}_{i}=\{s_{i,1},s_{i,2},...,s_{i,m}\}$. The overall uncertainty across all $n$ explanations is captured by aggregating reasoning-level uncertainties for each explanation. This can be formally defined as follows:
@@
Given an input prompt $x^{q}$ and an explanation-specific prompt $x^{e}$, the model $M$ generates a set of answers $a_{i}$ to the query $x^{q}$, along with accompanying explanations $a_{i}^{e}$. The uncertainty for the query $x^{q}$ is then defined as:
(3) $U_{x^{q}}=\mathcal{U}\left(\bigcup_{i=1}^{n}a^{e}_{i}\right)=\mathcal{U}\left(\bigcup_{i=1}^{n}\{s_{i,1},s_{i,2},...,s_{i,m}\}\right)$
@@
Here, $U_{x^{q}}$ represents the overall uncertainty for the query $x^{q}$, and $\mathcal{U}$ aggregates uncertainties across all reasoning steps from all $n$ explanations.
Unlike prior methods that focus on token-level or semantic variability (Tanneru et al., 2024), in this paper, we explicitly models reasoning structures within explanations. By leveraging logical topologies, we aim to capture nuanced uncertainties at both the explanation level and individual reasoning step level.
Figure 2. The illustration of how the topology is constructed.
4. Method
@@
In this section, we propose a novel framework for reasoning uncertainty quantification in large language models (LLMs). To capture the complexity of reasoning paths, the proposed framework consists of two main steps: (1) Reasoning Topology Elicitation, which constructs a structured reasoning graph from LLM-generated explanations, and (2) Topology-enabled Reasoning Quantification, which leverages the constructed graph to perform uncertainty quantification using measures such as graph edit distance and reasoning path redundancy analysis. These steps work together as a comprehensive framework for analyzing reasoning uncertainty in LLMs.
4.1. Reasoning Topology Elicitation
@@
The objective of this step is to construct a structured reasoning topology that captures the complexity of reasoning paths generated by large language models (LLMs). Existing approaches (Tanneru et al., 2024) represent reasoning explanation as linear sequences of steps from Chain-of-Thought (CoT) prompting (Wei et al., 2023). While linear text sequences provide basic interpretability, they fail to capture complex logical transitions required for tasks like comparative reasoning or multi-faceted conclusions. This lack of structural richness limits the ability to analyze and quantify uncertainty in LLM-generated reasoning processes.
@@
To address these limitations, we propose to elicit reasoning topologies from a question $x_{q}$ to an answer $a$ as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes corresponding to knowledge points or intermediate steps, and $\mathcal{E}$ is the set of edges capturing logical dependencies between them. For example, given the query $x^{q}$: “What are the causes of climate change?”, the knowledge points $K^{q}$ might include sub-questions such as $k_{1}$: “What is the role of greenhouse gases?”, $k_{2}$: “How does deforestation contribute?”, and $k_{3}$: “What is the impact of industrial activities?”. The corresponding answers $A=\{a_{1},a_{2},a_{3}\}$ provide detailed explanations for each sub-question towards the final answer $a_{3}$. These knowledge-answer pairs are then connected based on their logical dependencies to form the reasoning topology graph $\mathcal{G}^{q}$. This graph-based representation enables a more comprehensive understanding of the reasoning process and provides a foundation for richer uncertainty quantification.
@@
Specifically, the construction of $\mathcal{G}$ consists of three modules: (1) Knowledge Point Reflection, where the LLM identifies sub-questions or knowledge points required to address the query; (2) Self-Answering Module, where the LLM generates answers for each identified knowledge point; and (3) Reasoning Topology Construction, where the knowledge-answer pairs are organized into a directed graph that reflects the overall reasoning path.
4.1.1. Knowledge Point Reflection Module
Template $T_{1}$: Given a question {$x^{q}$}, reflect and generate the knowledge points or sub-questions necessary to solve this query. Ensure that the output is both sufficient and concise.
The model generates a set of knowledge points $K^{q}=M(x_{q},T_{1})=\{k_{1},k_{2},\dots,k_{n}\}$, where each $k_{i}$ corresponds to a specific sub-question or piece of information identified as necessary to address the query $x^{q}$ under the guidance of prompt $T_{1}$. To ensure traceability, we assign unique identifiers to each knowledge point using a tagging function $f(\cdot)$:
(4) $K^{q}_{\text{tag}}=\{\text{id}_{1}:k_{1},\text{id}_{2}:k_{2},\dots,\text{id}_{n}:k_{n}\}.$
In the later sections of this paper, we assume the $k^{q}$ always carries its identifier while performing computing ($K_{q}\Leftrightarrow K^{q}_{tag}$).
4.1.2. Self-answering Module
(5) $A=\{a_{1},a_{2},\dots,a_{n}\}=\{M\{k_{1},T_{2}\},M\{k_{2},T_{2}\},\dots,M\{k_{n},T_{2}\}\}$
(6) $\mathcal{D}_{m}=\{(k_{1},a_{1}),(k_{2},a_{2}),\dots,(k_{n},a_{n})\}$
4.1.3. Reasoning Topology Construction Module
(7) $\mathcal{\hat{D}}_{m}$ $=M(\mathcal{D}_{m},e^{F})=\{(a^{p_{1}},k_{1},a_{1}),(a^{p_{2}},k_{2},a_{2}),\dots,(a^{p_{n}},k_{n},a_{n})\}$
(8) $(a^{p_{1}},k_{1},a_{1})\Rightarrow(\underbrace{a^{p_{1}},a_{1}}_{{nodes}},\underbrace{k_{1}}_{edge})$
where the order of two nodes is defined by the reasoning LLM.
Now we can write a basic reasoning step as:
(9) $\mathcal{\text{Step}}_{ij}=[\text{node}_{i},\text{node}_{j},\text{edge}_{ij}]$
(10) $\mathcal{G}^{q}=(\mathcal{V},\mathcal{E}),$
where
(11) $\mathcal{V}=\{\texttt{nodeRaw},\text{node}_{1},\dots,\texttt{nodeResult}\}=\{v_{0},v_{1},...\}$
and the edges are expressed as:
(12) =$\mathcal{E}=\{e_{ij(1)},e_{ij(2)}...\}$
where $e$ stands for edge and $\{e_{ij}\mid\text{edge}_{ij}:\text{node}_{i}\to\text{node}_{j}\}$, $e_{ij}$ represents reasoning operations or dependencies between nodes. (start with index ‘1’ since we assume ‘0’ is the nodeRaw). The graph-based structure captures the full reasoning topology, including branching, dependencies, and multi-step interactions, which allows for better reflection of the relationships between intermediate steps.
Now from a graph concept, the reasoning steps combined with Eq. 9 are formalized as below:
(13) $S=\{\text{Step}_{ij}\mid\text{Step}_{ij}=[v_{i},v_{j},e_{ij}]v_{i},v_{j}\in\mathcal{V},e_{j}\in\mathcal{E}\}$
where each triplet represents a logical transition between reasoning steps. Note that for complex reasoning, the final answer does not necessarily rely on all of the reasoning steps. For example, when being asked about ”if currently is summer in Australia, what season is it in Canada?”, in the reasoning chain, some of the LLM might delve into ‘what causes the season differences’, which is redundant steps in a concise reasoning.
4.2. Topology-enabled Reasoning Quantification

### v2

@@
Abstract
It is important to understand the uncertainty in large language models (LLMs) explanations because they reveal more information about the reasoning process and thus provide insights into the reliability of LLMs’ answers regarding a question. In this work, we propose a novel framework that quantifies uncertainty in LLM explanations through reasoning topologies. By designing a structural elicitation strategy, we guide the LLM to frame the explanations on how it derives the answers into graph topologies. This strategy first queries knowledge-related sub-question and sub-answer pairs, and then guides the LLM to connect the pairs through a topological reasoning process. Based on the reasoning topologies, we revisit the Graph Edit Distance and provide a variant that can better quantify the LLM uncertainty from both semantic and reasoning structure dimensions. The topology structure further brings convenience to assess redundancy by extracting and comparing the valid reasoning path to the raw explanation. Extensive experiments show the effectiveness of the proposed framework, and interesting findings on reasoning patterns and efficiency are discussed.
1 Introduction
@@
Deep learning models have long been criticized for their lack of trustworthiness due to the complex network structures and opaque decision-making processes (Li et al., 2022; Doshi-Velez & Kim, 2017; Samek et al., 2021). This has motivated researchers to investigate methods for understanding and quantifying the uncertainty associated with these models (Abdar et al., 2021; Loquercio et al., 2020; Maddox et al., 2019). Recently, Large Language Models (LLMs) have demonstrated significant advancements over traditional deep learning approaches across a variety of tasks (Zhao et al., 2023; Naveed et al., 2023). However, concerns about their reliability persist. LLMs often produce outputs that are difficult to verify, particularly in scenarios requiring complex reasoning (Shi et al., 2024). This introduces risks in critical applications, such as healthcare or legal contexts (Cascella et al., 2023; Jayakumar et al., 2023), where incorrect or unreliable reasoning can have severe consequences. Properly quantifying uncertainty in reasoning is crucial for ensuring safe and trustworthy reference.
Existing research on Uncertainty Quantification (UQ) for LLMs primarily focuses on analyzing semantic uncertainty (Kuhn et al., 2023; Lin et al., 2023; Qiu & Miikkulainen, 2024; Da et al., 2024), which involves examining patterns from the meaning of multiple responses generated for a given question. Although this can provide insights into output-level variability, it overlooks the structural characteristics of the reasoning process that give rise to such uncertainty. In other words, existing UQ methods do not model how uncertainty emerges and propagates through intermediate reasoning steps, which hinders practitioners’ analyses, understanding, and potential remedy-based improvement. For instance, in Figure. 1 (right), when asked the same question, In part ① answer-based UQ can not reveal the vulnerable parts in the reasoning path, which are contained in the reasoning steps as in part ②.
Figure 1: The comparison of Answer-based-UQ measures and Proxy (Explanations)-based UQ measures. The Explanations-based approach reveals more reasoning information.
By quantifying uncertainty at the reasoning dimension, we can effectively detect inconsistencies in intermediate steps, support human-in-the-loop validation in sensitive applications, and uncover vulnerabilities within the reasoning process itself. This underscores the importance of reasoning-aware uncertainty quantification. In this paper, we address this challenge by explicitly modeling reasoning processes as logical topologies. While prior work treats a single Chain-of-Thought (CoT) sequence as a monolithic response (Wei et al., 2022; Wang et al., 2023) and measures semantic consistency at the output level, such approaches fail to capture the structural complexity of real-world reasoning, which frequently involves hierarchical dependencies and parallel sub-tasks. To overcome these limitations, we introduce formal topology that represents reasoning as a logical graph, where nodes correspond to individual reasoning steps and edges encode their logical dependencies. This graph-based structure enables finer-grained, more interpretable uncertainty analysis.
@@
  • We identify limitations in existing answer-based UQ approaches for LLMs and propose a novel topology-elicitation framework that explicitly models the reasoning process from explanations as a logical topology.
  • We introduce Topo-UQ, a topology-based uncertainty measure from both semantic meaning and reasoning structure to show granular thinking variance. We also propose a redundancy metric based on valid reasoning paths and raw explanations.
  • We demonstrate the effectiveness of our framework through extensive experiments across multiple datasets and LLMs. Our results highlight its ability to identify inconsistencies in reasoning paths, improving trustworthiness and interpretability. Additionally, we discover three widely adopted reasoning patterns in LLMs and show the chance of improvement under the redundancy measure.
2 Related Work
@@ 2.1 UQ for LLM
A significant body of research has focused on performing UQ for LLMs by inducing the models to output their uncertainty along with their responses Kadavath et al. (2022); Lin et al. (2022); Mielke et al. (2020); Tian et al. (2023). These methods often rely on token-level probabilities to train or fine-tune models for predicting uncertainty. While effective, these approaches require full access to the model’s structure and weights, which is impractical for black-box or commercial LLMs. For example, supervised methods such as those in Kadavath et al. (2022) estimate uncertainty using logits and ground truth labels but are computationally expensive and resource-intensive.
@@
Another line of work estimates uncertainty directly at the response level using semantic entropy Kuhn et al. (2023). While this method avoids token-level dependencies, it still relies on access to token probabilities, limiting its applicability in black-box settings. To address these limitations, researchers have proposed lightweight black-box methods that analyze response inconsistencies. For instance,  Lin et al. (2023) uses graph Laplacian eigenvalues as an uncertainty indicator, while  Chen & Mueller (2023) computes confidence scores from generated outputs to identify speculative or unreliable answers. However, these approaches primarily focus on semantic-level analysis and neglect the logical structure underlying reasoning processes. Moreover, methods like Lin et al. (2023) average entailment probabilities without considering directional information in reasoning paths.
@@
Our work is agnostic to the white box or black box since it leverages the generated explanations as a proxy to measure the reasoning uncertainty as in Figure. 1. This enables a more nuanced and interpretable assessment of uncertainty in reasoning processes.
2.2 UQ for LLM Explanation
@@
Explanation-based UQ focuses on assessing the reliability of natural language explanations (NLEs) generated by LLMs by either prompting models to express confidence in their explanations or analyzing consistency across multiple outputs under varying conditions Tanneru et al. (2024); Yadkori et al. (2024). While these methods provide insights into explanation robustness, they treat explanations to a question as unstructured text representation, which lacks structural information and fails to capture inconsistencies or leaps in logic. In contrast, our work explicitly leverages well-structured reasoning topologies to enhance the UQ process for explanations. This structured representation enables us to assess explanation uncertainties at a finer granularity within complex reasoning paths.
3 Preliminaries
@@
This section provides the foundations for our study, including the definition of uncertainty in LLMs and its quantification in natural language explanations (NLE) fashion.
3.1 Uncertainty Quantification of LLMs
Uncertainty, as an inherent attribute of LLMs, can arise from various aspects: training corpus coverage (Huang et al., 2023), prompt ambiguity (Abbasi Yadkori et al., 2024), the probabilistic nature of language generation (Lin et al., 2023), and many more. Effectively modeling and interpreting uncertainty are essential for reliable and trustworthy references; thus, researchers quantify the uncertainty for better decision-making.
Different from UQ methods in classical machine learning (Lakshminarayanan et al., 2017; Gal & Ghahramani, 2016; Hernández-Lobato & Adams, 2015; Abdar et al., 2021), UQ in LLM faces unique challenges such as discrete token-based generation, the large combinatorial output space, and expression diversities of the same concept, thus, most LLM-UQ methods focus on analyzing uncertainty from model-generated responses as shown in ① in Figure. 1. The problem could be defined as follows:
@@ 3.1 Uncertainty Quantification of LLMs
The LLM ‘$M$’ is provided with an input query $x^{q}$, the goal of the uncertainty function $U_{x}$ is to map the generated outputs to a scalar score that determines the uncertainty of the $M$, i.e., $U_{x}=\mathcal{U}\left(\{M(x^{q}_{i})\}_{i=1}^{n}\right)$
Here, $\{M(x^{q}_{i})\}_{i=1}^{n}$ denotes a set of $n$ responses [^1] Answers and responses are used interchangeably in this paper. generated by the model $M$, and $\mathcal{U}$ aggregates uncertainty across multiple responses. $U_{x}$ is calculated only on $x$ and is a property of the predicted distribution, which is estimated by $\mathcal{U}$. For black-box methods that analyze response variability or semantic consistency (Lin et al., 2023), multiple outputs ($n>1$) are typically required. In contrast, white-box methods that rely on internal model parameters, such as logits (Kadavath et al., 2022), may only require a single output ($n=1$). However, both current answer-based black-box Lin et al. (2023); Chen et al. (2025); Su et al. (2024) and white-box methods Margatina et al. (2023); Duan et al. (2024); Kadavath et al. (2022) share limitations in the quantification of reasoning procedures in that they focus on semantic variability without capturing deeper uncertainties in reasoning steps.
3.2 Uncertainty of LLM Explanations
To address these issues, researchers try to understand uncertainties in reasoning processes through the natural language explanations (NLEs) as a proxy, introduced by  Camburu et al. (2018); Tanneru et al. (2024). As shown in the green part (②) of Figure. 1, an NLE is a textual reasoning sequence generated by a language model $M$, to justify or explain the derivation of answer $a$ for a given input question $x^{q}$. Following the existing work, it is defined as:
@@ 3.2 Uncertainty of LLM Explanations
Given a model $M$ and an input query $x^{q}$, an NLE can be represented as:
$M(x^{q}+x^{e})=a+a^{e}$ (1)
where $x^{e}$ is an explanation-specific prompt, $a$ is the model’s answer to the query $x^{q}$, and $a^{e}$ is the generated explanation accompanying the answer. The explanation $a^{e}$ contains a sequence of reasoning steps, represented as $a^{e}=\{s_{1},s_{2},...,s_{m}\}$, which capture the reasoning process or justification for $a$.
To quantify uncertainty through explanations, we extend Problem 1 to consider $n$ explanations generated for the same query $x^{q}$. Each explanation $a^{e}_{i}$ ($i\in\{1,2,\dots,n\}$) corresponds to a set of reasoning steps, assuming $m$ steps, then we denote them as $a^{e}_{i}=\{s_{i,1},s_{i,2},...,s_{i,m}\}$. The overall uncertainty across all $n$ explanations is captured by aggregating reasoning-level uncertainties for each explanation. This can be formally defined as:
@@
Given an input query $x^{q}$ and an explanation-specific prompt $x^{e}$, the model $M$ generates a set of answers $a_{i}$ to the query $x^{q}$, along with accompanying explanations $a_{i}^{e}$. The uncertainty for the query $x^{q}$ is then defined as:
$U_{x^{q}}=\mathcal{U}\left(\bigcup_{i=1}^{n}a^{e}_{i}\right)=\mathcal{U}\left(\bigcup_{i=1}^{n}\{s_{i,1},s_{i,2},...,s_{i,m}\}\right)$ (2)
@@
where $U_{x^{q}}$ represents the overall uncertainty for the query $x^{q}$, and $\mathcal{U}$ aggregates uncertainties across all reasoning steps from all $n$ explanations.
Unlike prior methods that focus on answer-based output variability (Tanneru et al., 2024), in this paper, we aim to capture nuanced uncertainties from the reasoning process, which brings advantages in analyzing the reasoning patterns and logical errors. In the next section, we will introduce Topo-UQ, including how to model reasoning structures within explanations, and how this graph-based UQ method works with the extracted structures.
Figure 2: An overview of Topo-UQ. It depends on two modules, ① Topology Elicitation and ② Uncertainty Quantification. Start from ①, given an input question $x^{q}$, the LLM not only generates the answer $a$ but is also elicited to provide reasoning explanations. It first reflects the necessary knowledge points ($k_{n}$) as sub-questions and then answers them as ($a_{n}$), $T_{1}$ and $T_{2}$ are the prompt templates. These sub-question and answer sets form the edge and node sets $\mathcal{E}$, $\mathcal{V}$, creating matched pairs. By inserting a node into an (edge-node) pair, the LLM constructs the reasoning topology $\mathcal{G}$. When we have $L$ explanations, their topologies are used to estimate Uncertainty through module ②: we encode the content from each node and edge into embedding space to capture semantic meanings, and apply a variant of Graph Edit Distance to measure structural variance, then the final uncertainty combines both semantic and structural differences. An example is shown at the bottom.
4 Method
@@
In this section, we propose a reasoning Topology-based method for Uncertainty Quantification of LLMs, named Topo-UQ. To capture the complexity of reasoning paths, the proposed framework consists of two modules: (1) Reasoning Topology Elicitation, which constructs a structured reasoning graph from LLM-generated explanations, and (2) Topology-based UQ, which leverages the constructed graphs for uncertainty quantification using a variant of graph edit distance. We also present a byproduct - a metric of reasoning path redundancy inspired by the analysis of reasonable reasoning paths and original explanations.
4.1 Reasoning Topology Elicitation
@@
The objective of this step is to derive a structured reasoning topology that captures the complexity of reasoning paths generated by large language models (LLMs). Existing approaches by Tanneru et al. (2024) represent reasoning explanation as linear sequences of steps from Chain-of-Thought (CoT) (Wei et al., 2023). While linear text sequences provide basic interpretability, they fail to capture complex logical transitions in tasks like comparative reasoning or multi-faceted conclusive answering. This lack of structural richness limits their ability to analyze and quantify uncertainty in LLM-generated reasoning processes.
@@
To address these limitations, we propose to elicit reasoning topologies from a question $x_{q}$ to an answer $a$ as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes corresponding to knowledge points or intermediate steps, and $\mathcal{E}$ is the set of edges capturing logical dependencies between them. The $\mathcal{G}$ includes two advantages: First, it can represent diverse structures, not limited to the linear sequence CoT. Second, the construction is done by the LLM itself, thus it can flexibly reflect the thinking process for each answer. For example in Figure. 2, given the query $x^{q}$: “If it is currently Summer in Australia, what season is it in Canada?”, the knowledge points might include sub-questions such as $k_{0}$: ‘Where is Australia located on Earth?’, $k_{1}$: ‘Where is Canada located on Earth?’, $k_{2}$: ‘What is the geographical relationship between Australia and Canada?’ and $k_{3}$: ‘How does the tilt of the Earth affect seasons?’. The corresponding answers $A=\{a_{0},a_{1},a_{2},a_{3}\}$ demonstrate thoughts of LLM for sub-questions. These knowledge-answer pairs are then connected based on their logical dependencies to form the reasoning topology graph $\mathcal{G}$. The connection becomes a step from $a_{i}$ to $a_{j}$: $S=\{[a_{i},a_{j},k_{ij}]\mid a_{i},a_{j}\in\mathcal{V},k_{ij}\in\mathcal{E}\}$
@@
Specifically, the construction of $\mathcal{G}$ consists of three steps: (1) Knowledge Point Reflection, where the LLM identifies sub-questions of knowledge points $k_{n}$ required to address the query as elicited by the template $T_{1}$; (2) Self-Answering Module, based on the sub-questions among $k_{n}$, the LLM generates answers for them, denoted as $a_{n}$, in this process, template $T_{2}$ carries the question variable which can be changed according to different sub-questions; up to this point, there is an intermediate set containing pairs222($k_{n}$, $a_{n}$) = (edge, node) = ($e$, $v$) ($k_{n}$, $a_{n}$), which will be used for (3) Reasoning Topology Construction. In this part, the knowledge-answer pairs are connected into a directed graph that reflects the overall reasoning process. It mainly requires an insertion from the current node-edge pair to a prior node, which is identified by LLM itself. The detailed explanations of every module and prompt templates are introduced in the Appendix. B.
4.2 Topology-based Uncertainty Quantification